# TopoCode: Topologically Informed Error Detection and Correction in Communication Systems

### v2

@@ Abstract
Traditional error detection and correction codes focus on bit-level fidelity, which is insufficient for emerging technologies like eXtended Reality (XR) and holographic communications requiring high-data-rate, low-latency systems. Bit-level metrics cannot comprehensively evaluate Quality-of-Service (QoS) in these scenarios. This letter proposes TopoCode which leverages Topological Data Analysis (TDA) and persistent homology to encode topological information for message-level error detection and correction. It introduces minimal redundancy while enabling effective data reconstruction, especially in low Signal-to-Noise Ratio (SNR) conditions. TopoCode offers a promising approach to meet the demands of next-generation communication systems prioritizing semantic accuracy and message-level integrity.
@@ I Introduction
Channel coding for error detection and correction is essential in data communication to ensure reliability and achieve high effective data rates. Various codes, such as linear block codes, cyclic redundancy check codes, and convolutional codes, have been extensively developed and applied [1]. These codes primarily address binary source data and enable high-fidelity decoding even in the presence of channel distortion and noise. While accurate binary decoding is desirable, it is important to note that not all bits are equally significant, and certain errors may have minimal impact post-decoding. For instance, in eXtended Reality (XR) and Holographic-Type Communication (HTC)[2, 3], point cloud data with large sizes is common. Communication errors may result in slight shifts in point locations; however, if the overall geometry remains unaffected, the errors are often imperceptible. Similarly, in 360-degree videos, errors at the boundary of a user’s Field-of-View that do not alter the image’s geometry are negligible.
@@
Semantic communication encodes source data by leveraging domain knowledge and its semantic meaning [4]. Joint source and channel coding in semantic communication utilizes deep learning models to extract and encode semantic information effectively. At the receiver side, this highly compressed semantic information is decoded using similar models. While semantic communication efficiently encodes source data, it relies heavily on specialized deep learning models. Training a unified model for arbitrary data transmission remains a significant challenge. Moreover, deep learning requires extensive training data, which is not always available, and the trained models often lack generalization. Thus, there is a need for a novel approach that avoids training, applies to any data modality, and focuses on detecting and correcting critical errors without addressing all errors indiscriminately.
@@
In this letter, we propose TopoCode, a novel approach leveraging Topological Data Analysis (TDA) [5, 6], particularly persistent homology and persistence diagrams, to extract and represent data’s topological information. This information focuses on the shape and geometry of the data, capturing high-level relations and features. Similar to semantic meaning, meaningful data often exhibits underlying manifold structures that can be extracted through TDA. In TopoCode, the encoded persistence diagram is concatenated with the source data and transmitted through the communication channel. At the receiver, the decoded data is used to generate a new persistence diagram, which is compared with the received one. This allows for error detection in the topological space, evaluation of error significance, and correction of topological errors through optimization. The advantages of TopoCode are as follows. First, TopoCode introduces minimal redundancy to the source data while effectively detecting errors and assessing their significance using topological information. Second, it can correct topological errors and recover the source data, particularly in low Signal-to-Noise Ratio (SNR) scenarios. Third, TopoCode utilizes established computational topology tools, eliminating the need for training or pre-collected data. This enables its application across various data modalities, including video, point cloud, image, and time-series data, which has broad generalizability.
@@ II Topological Data Analysis
Topological Data Analysis (TDA) is used to study the topological features across various data modalities. It identifies different homology groups, such as the $0^{\text{th}}$ group ($H_{0}$), representing connected components, the $1^{\text{st}}$ group ($H_{1}$), representing holes or loops, and the $2^{\text{nd}}$ group ($H_{2}$), representing voids. TDA measures the size of each homology group, providing insights into the overall shape and structure of the data. However, size alone cannot capture the evolution of these homology groups or their relative relationships. To address this limitation, persistent homology is employed to track the evolution of each component in a homology group through Birth-Death (BD) pairs, capturing richer topological information. For example, a point cloud circle and its corresponding persistence diagram are shown on the left of Fig. 1. In the $H_{0}$ homology group, BD pairs are born at 0 and die at various values because each point starts as an individual connected component. The birth and death values represent the radii of disks centered at the points. As these disks grow, when two disks touch, one component dies, while the other persists. In the $H_{1}$ homology group, a BD pair $(0.66,1.76)$ represents the circular structure. The circle is formed when the radii of the disks reach approximately 0.66, and as the radii increase to 1.76, the circular structure closes. This process is called the Vietoris-Rips Filtration, where the radius serves as the filtration value [5, 6].
@@
(d) Persistence diagram of the circle.
(e) Persistence diagram of noisy circle.
(f) Persistence diagram of 3/4 circle.
Figure 1: Comparison of point clouds and persistence diagrams for circle variants.
In the middle and on the right-hand side of Fig. 1, we show two circular point clouds which are distorted mildly and severely, respectively. This can be reflected in the persistence diagram. First, for the middle point cloud, the $H_{1}$ homology group BD pair shifts more towards the diagonal which locates at (1.06,1.72). Since death must be after birth, all BD pairs are above the diagonal. However, if a BD pair is close to the diagonal, it means the BD pair quickly dies after its birth which represents insignificant topological feature. Thus, the circular structure is weaker compared with the one in the left figure. Second, for the right point cloud, the $H_{1}$ homology group is almost on the diagonal which locates at (1.69,1.77). This shows that the circular structure is not strong which can be neglected, which indicates a significant difference from the one in the left figure.
For more complex figures, the size of homology group becomes large and unknown. Since persistent homology captures the intrinsic topological information in the data, the size of the homology group depends on the data which is not a constant. When the number of BD pairs becomes large, it is challenging to evaluate the change of persistence diagram. In persistent homology, there are well-defined non-Euclidean distances. Consider that the source data is ${\bm{X}}$ and the distorted data is ${\hat{\bm{X}}}$. Then, we perform filtration using a function $f_{l}(\cdot)$ (such as Vietoris-Rips Complex) and obtain persistence diagram $D(f_{l}(\cdot))$. The difference between ${\bm{X}}$ and ${\hat{\bm{X}}}$ can be evaluated using their persistence diagrams [7],
@@
where $\Delta$ is the diagonal of the persistence diagram, $\|(a,b)\|_{\infty}=\max\{|a|,|b|\}$ and $\eta(x)$ is a bijection function which maps the BD pairs in $D(f_{l}({\bm{X}}))$ to those in $D(f_{l}(\hat{\bm{X}}))$, i.e., $\eta:D(f_{l}({\bm{X}}))\cup\Delta\rightarrow D(f_{l}(\hat{\bm{X}}))\cup\Delta$. Since the two diagrams may have different number of BD pairs, some of the BD paris that cannot be matched are mapped to the diagonal $\Delta$. The above distance measurement is called $p$-Wasserstein distance.
Figure 2: Wassertein distance between the examples in Fig. 1 with different $p$ values.
@@
An example of the $p$-Wasserstein distance between the point clouds in Fig. 1 is shown in Fig. 2. The $p$-Wasserstein distance includes the summation of the $H_{0}$ and $H_{1}$ distances. The noisy circle and the 3/4 circle is compared with the original circle. As we can see, the noisy circle has a smaller $p$-Wasserstein distance when $p$ is from 1 to 5. In this paper, without specific notation, we use $p=2$, which is a widely used parameter. Note that, the Wasserstein distance between two identical diagrams is 0. Thus, the distance can be used to evaluate the received data’s topological fidelity.
@@ III Error Detection and Correction
In this section, we introduce the TopoCode structure and the algorithms for error detection and correction using persistence diagrams. An illustration of the TopoCode data packet structure is shown in Fig. 3. The source data, which can be images, time series data, and point clouds, are uncoded. The packet meta data can contain information about the source data length. The TopoCode is concatenated with the source data and it is organized by the order of homology groups. Each homology group contains its length and the homology group data. Although only $H_{0}$ and $H_{1}$ are shown, the homology group number can be larger depending on the application, e.g., 3D point cloud data may use $H_{2}$.
The homology group data, e.g., $H_{0}$ and $H_{1}$ data, can have many different formats. First, the persistence diagram can be directly encoded using the coordinates of BD pairs. The encoded homology group data is a series of BD pairs. However, the number of BD pairs is unknown which results in variable code length. To address this issue, there are various other fixed-length topological features that can be derived from persistence diagrams, such as the persistence landscape, persistence image, quantized persistence diagram, and amplitude-based features [6]. Although they can generate fixed-length code, we lose complete information of the original persistence diagram and topological distance metrics such as Wasserstein distance cannot be used. Since this is the first paper uses topological information for error detection and correction, we use the original persistence diagram. Other fixed-length or quantized topological features will be studied in future works.
Each data modality can be processed by specific filtration approaches due to its unique format. For example, Cubical Complex is more efficient for images due to their grid structures, and Alpha Complex and Vietoris-Rips Complex are suitable for Point Clouds since the points are distributed in 3D space without any specific format. Due to the limited space, this letter only considers gray images and uses Cubical Complex to obtain the persistence diagrams [8].
In persistent homology, the stability of persistence diagram for grayscale functions is ensured in the following way [9]:
@@
where $C_{fil}$ is a coefficient which depends on the filtration types and the data values. The above inequality shows that $C_{fil}$ times the difference of two gray images is no smaller than their $p$-Wasserstein distance. If two images are similar, the $p$-Wasserstein distance must be small. As a result, the topological errors in the received data can be detected using persistence diagrams.
@@
Equation (2) provides the distance between ${\bm{X}}$ and ${\hat{\bm{X}}}$. If we change ${\hat{\bm{X}}}$ to noises, Eq. (2) shows the topological information in ${\bm{X}}$, i.e., the distance to noises in the topological space. As we have discussed, the noises in persistence diagrams mainly appear around the diagonal, which can be considered as the noise floor. As shown in Fig. 4, a persistence diagram contains three BD pairs with distance to diagonal being $d_{1}$, $d_{2}$, and $d_{3}$. If we increase the noise floor from $p_{0}$ to $p_{1}$, the BD pair 3 is considered as noises which is not important and can be removed. Then, we obtain an updated denoised persistence diagram with BD pairs 1 and 2. We can further increase the noise floor to $p_{2}$ and only keep significant BD pairs. In order to select the $p_{1}$ and $p_{2}$, we use the total persistence in this letter, which is the summation of all the finite distances of BD pairs to the diagonal
$T_{t}^{h}=\sum_{i\in D_{h}(f_{l}({\bm{X}}))}d_{i},{\text{ for }}h\in\{0,1,\cdots\},$ (3)
where $D_{h}(f_{l}({\bm{X}}))$ represents the $h$-th homology group in the persistence diagram. In this case, we consider the homology groups individually. Each homology group $h$ has its own total persistence $T_{t}^{h}$. The total persistence can represent the strength of the topological feature. For instance, Gaussian noises have nearly 0 total persistence.
@@
For error detection, we compare the received data’s persistence diagram with TopoCode. However, noises can corrupt the transmitted data and the Wasserstein distance in Eq. (2) can be large. In order to reduce the impact of noises, we remove the insignificant BD pairs which can be introduced by noises by choosing a threshold percentage $\alpha\in[0,1]$ and set new noise floor $p^{h}=\alpha T_{t}^{h}$. All BD pairs below $p^{h}$ are removed and we only compare the significant BD pairs. If removing insignificant BD pairs and the distance between the updated persistence diagram is small to the original one, the significant topological feature still remains in the received data. The errors can be accepted without requesting for retransmission. The selection of $\alpha$ depends on how much error can be tolerated. We can gradually increase $\alpha$ to evaluate the similarity at different persistence levels. The $1-\alpha$ can be considered as the overall significance level. The larger $\alpha$, the smaller overall significance.
@@
Finally, we can obtain the Wasserstein distance at different significance level which represents the detected error significance. For example, if the Wasserstein distance decreases as $\alpha$ increase, that means most of the errors are insignificant and the major topological feature of the data remains. On the contrary, if the Wasserstein distance does not change as $\alpha$ increase, that means the error is significant which may require retransmission. Note that, the TopoCode error detection evaluate the data at the message topology level. The received data can have insignificant bit-level errors but TopoCode cannot detect those.
@@
Given the received data and the TopoCode, we can detect their mismatch using the error detection. In addition, we can correct the received data with topological errors using the received TopoCode in the following way:
@@
where $H$ is number of homology groups used, ${\mathcal{D}}$ is the set where the source data ${\bm{X}}$ is defined. The idea is to minimize the $p$-Wasserstein distance and obtain the recovered source data ${\tilde{\bm{X}}}$.

### v3

# Topocode: Topologically Informed Error Detection and Correction in Image Communication Systems

@@ Abstract
Immersive communication technologies such as eXtended Reality (XR) and holographic-type communications require high-data-rate low-latency services. Due to the massive amount of transmitted data, bit-level metrics cannot comprehensively evaluate the Quality-of-Experience (QoE) in these scenarios. This letter proposes Topocode which leverages Topological Data Analysis (TDA) and persistent homology to encode topological information for message-level error detection and correction in image communication systems. It introduces minimal redundancy while enabling effective data reconstruction, especially in low Signal-to-Noise Ratio (SNR) scenarios. Topocode offers a promising approach to meet the demands of next-generation communication systems prioritizing message-level integrity.
@@ I Introduction
Channel coding for error detection and correction is essential in data communication to ensure reliability and achieve high effective data rates. Various codes, such as linear block codes, cyclic redundancy check codes, and convolutional codes, have been extensively developed and applied [1]. These codes primarily address binary source data and enable high-fidelity decoding even in the presence of channel distortion and noise. While accurate binary decoding is desirable, it is important to note that not all bits are equally significant, and certain errors may have minimal impact post-decoding. For instance, in eXtended Reality (XR) and Holographic-Type Communication (HTC)[2, 3], point cloud data with large sizes is common. Communication errors may result in slight shifts in point locations; however, if the overall geometry remains unaffected, the errors are often imperceptible.
@@
Semantic communication encodes source data by leveraging domain knowledge and its semantic meaning [4]. Joint source and channel coding in semantic communication utilizes Deep Learning (DL) models to extract and encode semantic information effectively. At the receiver side, this highly compressed semantic information is decoded using DL decoding models. While semantic communication efficiently encodes source data, it utilizes application-specific DL models. Training a unified model for arbitrary data transmission remains a significant challenge. Moreover, DL requires extensive training data, which is not always available, and the trained models often lack generalization. Thus, there is a need for a novel approach that avoids training, applies to any data modality, and focuses on detecting and correcting critical errors without addressing all errors indiscriminately.
@@
In this letter, we propose Topocode, a novel approach leveraging Topological Data Analysis (TDA) [5, 6], particularly persistent homology and Persistence Diagram (PD), to extract and represent image data’s topological information. This approach has not been explored in existing coding schemes. TDA focuses on the shape and geometry of the data, capturing high-level relations and features. Similar to semantic meaning, meaningful data often exhibits underlying manifold structures that can be extracted through TDA. In Topocode, the encoded PD is concatenated with the source data and transmitted through the communication channel. At the receiver, the decoded data is used to generate a new PD, which is compared with the received one. This allows for error detection in the topological space, evaluation of error significance, and correction of topological errors through optimization. The advantages of Topocode are as follows. First, Topocode introduces minimal redundancy to the source data while effectively detecting message-level errors and assessing their significance. Second, it can correct topological errors, particularly in low Signal-to-Noise Ratio (SNR) scenarios. Third, Topocode utilizes established Computational Topology tools which eliminates the need for training or pre-collected data. This enables its application across various data modalities besides images, such as point cloud data, time-series data, and text, which has broad generalizability.
@@ II Topological Data Analysis
TDA is used to study the topological features across various data modalities. It identifies different homology groups, such as the $0^{\text{th}}$ group ($H_{0}$), representing connected components, the $1^{\text{st}}$ group ($H_{1}$), representing holes or loops, and the $2^{\text{nd}}$ group ($H_{2}$), representing voids. TDA measures the size of each homology group, providing insights into the overall shape and structure of the data. However, size alone cannot capture the evolution of these homology groups or their relative relationships. To address this limitation, persistent homology is employed to track the evolution of each component in a homology group through Birth-Death (BD) pairs which can capture richer topological information. For example, a point cloud circle and its corresponding PD are shown on the left side of Fig. 1. In the $H_{0}$ homology group, BD pairs are born at 0 and die at various values because each point starts as an individual connected component. The birth and death values represent the radii of disks centered at the points. As these disks grow, when two disks touch, one component dies, while the other persists. In the $H_{1}$ homology group, a BD pair $(0.66,1.76)$ represents the circular structure. The circle is formed when the radii of the disks reach approximately 0.66, and as the radii increase to 1.76, the circular structure closes. This process is called the Vietoris-Rips Filtration, where the radius serves as the filtration value [5, 6].
@@
(d) PD of the circle.
(e) PD of noisy circle.
(f) PD of 3/4 circle.
Figure 1: Comparison of point clouds and PDs for circle variants.
In the middle and on the right-hand side of Fig. 1, we show two circular point clouds which are distorted mildly and severely, respectively. This can be reflected in the PD. First, in Fig. 1(e), the $H_{1}$ homology group BD pair shifts more towards the diagonal which locates at (1.06,1.72). Since death must be after birth, all BD pairs are above the diagonal. However, if a BD pair is close to the diagonal, it means the BD pair quickly dies after its birth which represents insignificant topological feature. Thus, the circular structure is weaker compared with the one in Fig. 1(a). Second, for the point cloud in Fig. 1(c), the $H_{1}$ homology group is almost on the diagonal which locates at (1.69,1.77) in Fig. 1(f). This shows that the circular structure is not strong which can be neglected, which indicates a significant difference from the one in Fig. 1(a).
Since persistent homology captures the intrinsic topological information in the data, the size of the homology group depends on the data which is not a constant. When the number of BD pairs becomes large, it is challenging to evaluate the change of PD. In persistent homology, there are well-defined non-Euclidean distances. Consider that the source data is ${\bm{X}}$ and the distorted data is ${\hat{\bm{X}}}$. Then, we perform filtration using a function $f_{l}(\cdot)$ (such as Vietoris-Rips Complex) and obtain PD $D(f_{l}(\cdot))$. The difference between ${\bm{X}}$ and ${\hat{\bm{X}}}$ can be evaluated using their PDs [7],
@@
where $\Delta$ is the diagonal of the PD, $\|(a,b)\|_{\infty}=\max\{|a|,|b|\}$ and $\eta(x)$ is a bijection function which maps the BD pairs in $D(f_{l}({\bm{X}}))$ to those in $D(f_{l}(\hat{\bm{X}}))$, i.e., $\eta:D(f_{l}({\bm{X}}))\cup\Delta\rightarrow D(f_{l}(\hat{\bm{X}}))\cup\Delta$. Since the two diagrams may have different number of BD pairs, some of the BD paris that cannot be matched are mapped to the diagonal $\Delta$. The above distance measurement is called $p$-Wasserstein distance.
Figure 2: Wasserstein distance between the examples in Fig. 1 with different $p$ values.
@@
An example of the $p$-Wasserstein distance between the point clouds in Fig. 1 is shown in Fig. 2. The $p$-Wasserstein distance includes the summation of the $H_{0}$ and $H_{1}$ distances. The noisy circle and the 3/4 circle are compared with the original circle. As we can see, the noisy circle has a smaller $p$-Wasserstein distance when $p$ is from 1 to 5. In this paper, without specific notation, we use $p=2$, which is a widely used parameter. Note that, the Wasserstein distance between two identical diagrams is 0. Thus, the distance can be used to evaluate the received data’s topological fidelity.
@@ III Error Detection and Correction
In this section, we introduce the Topocode structure and the algorithms for error detection and correction using PDs. An illustration of the encoded Topocode structure is shown in Fig. 3. The source data, which is an image or a segment of image, is uncoded. The packet meta data can contain information about the source data length. Topocode is concatenated with the source data and it is organized by the order of homology groups. Each homology group contains its length and the homology group data. Although only $H_{0}$ and $H_{1}$ are shown, the homology group number can be varied depending on the application, e.g., 3D point cloud data may use $H_{2}$.
The homology group data, e.g., $H_{0}$ and $H_{1}$ data, can have many different formats. First, the PD can be directly encoded using the coordinates of BD pairs. The encoded homology group data is a series of BD pairs. However, the number of BD pairs is unknown which results in variable code length. To address this issue, there are various other fixed-length topological features that can be derived from PDs, such as the persistence landscape, persistence image, quantized PD, and amplitude-based features [6]. Although they can generate fixed-length code, we lose complete information of the original PD and topological distance metrics such as Wasserstein distance cannot be used. Since this is the first paper uses topological information for error detection and correction, we use the original PD which contains all available topological information. Other fixed-length or quantized topological features will be studied in future works.
Different data modalities can be processed by specific filtration approaches due to unique data formats. For example, Cubical Complex is more efficient for images due to their grid structures, and Alpha Complex and Vietoris-Rips Complex are suitable for Point Clouds since the points are distributed in 3D space without any specific format. This letter only considers gray images and uses Cubical Complex to obtain the PDs [8].
In persistent homology, the stability of PD for grayscale functions is ensured in the following way [9]:
@@
where $C_{fil}$ is a coefficient which depends on the filtration types and the data values. The above inequality shows that $C_{fil}$ times the difference of two gray images is no smaller than their $p$-Wasserstein distance. If two images are similar, the $p$-Wasserstein distance must be small. As a result, the topological errors in the received data can be detected using PDs.
@@
Equation (2) provides a lower bound of the distance between ${\bm{X}}$ and ${\hat{\bm{X}}}$. If we change ${\hat{\bm{X}}}$ to noises, Eq. (2) shows the topological information in ${\bm{X}}$, i.e., the distance to noises in the topological space. As we have discussed, the noises in PDs mainly appear around the diagonal, which can be considered as the noise floor. As shown in Fig. 4, a PD contains three BD pairs with distances to diagonal being $d_{1}$, $d_{2}$, and $d_{3}$. If we increase the noise floor from $p_{0}$ to $p_{1}$, the BD pair 3 is considered as noises which is not important and can be removed. Then, we obtain an updated denoised PD with BD pairs 1 and 2. We can further increase the noise floor to $p_{2}$ and only keep significant BD pairs. In order to select the $p_{1}$ and $p_{2}$, we use the total persistence in this letter, which is the summation of all the finite distances of BD pairs to the diagonal $T_{t}^{h}=\sum_{i\in D_{h}(f_{l}({\bm{X}}))}d_{i},{\text{ for }}h\in\{0,1,\cdots\},$ where $D_{h}(f_{l}({\bm{X}}))$ represents the $h$-th homology group in the PD. In this case, we consider the homology groups individually. Each homology group $h$ has its own total persistence $T_{t}^{h}$. The total persistence can represent the strength of the topological feature. For instance, Gaussian noises with limited number of BD pairs have nearly 0 total persistence.
@@
For error detection, we compare the received data’s PD with Topocode. However, noises can corrupt the transmitted data, and the Wasserstein distance in Eq. (2) can be large. In order to reduce the impact of noises, we remove the insignificant BD pairs, which can be introduced by noises, by choosing a threshold percentage $\alpha\in[0,1]$ and set new noise floor $p^{h}=\alpha T_{t}^{h}$. All BD pairs below $p^{h}$ are removed and we only compare the significant BD pairs. If removing insignificant BD pairs and the distance between the updated PD is small to the original one, the significant topological feature still remains in the received data. The errors can be accepted without requesting for retransmission. The selection of $\alpha$ depends on how much error can be tolerated. We can gradually increase $\alpha$ to evaluate the similarity at different persistence levels. The $1-\alpha$ can be considered as the overall significance level. The larger $\alpha$, the smaller overall significance.
@@
Finally, we can obtain the Wasserstein distance at different significance level which represents the detected error significance. For example, if the Wasserstein distance decreases as $\alpha$ increase, that means most of the errors are insignificant and the major topological feature of the data remains. On the contrary, if the Wasserstein distance does not change much as $\alpha$ increase, that means the error is significant which may require retransmission. Note that, Topocode error detection evaluate the data at the message topology level. The received data can have insignificant bit-level errors but Topocode cannot detect those.
@@
Given the received data and Topocode, we can detect their mismatch using the error detection. In addition, we can correct the received data with topological errors using the received Topocode in the following way:
@@
where $H$ is number of homology groups used, ${\mathcal{D}}$ is the set where the source data ${\bm{X}}$ is defined. The idea is to minimize the $p$-Wasserstein distance and obtain the recovered source data ${\tilde{\bm{X}}}$. In Eq. 3, the $D_{h}(f_{l}({\bm{X}}))$ is the Topocode sent by the transmitter.
In order to solve Eq. (3), we can use gradient decent. It is essential to ensure that the filtration and the $p$-Wasserstein distance computation are differentiable. The optimization of persistent homology functions has been studied in [10]. The PD is formed by BD pairs which are generated by complexes. The complex is generated based on distance functions which are differentiable. As a result, we can solve Eq. (3) using the following steps. First, ${\hat{\bm{X}}}$ is normalized by dividing their maximum values before obtaining the PD, so as the ${{\bm{X}}}$ at the transmitter side before generating Topocode. Second, given the target PD $D(f_{l}({\bm{X}}))$, we define the optimization loss function, which is
$\sum_{i}\min(|{\hat{\bm{X}}}_{[i]}|,|1-{\hat{\bm{X}}}_{[i]}|),$ (4)
The complexity to obtain the persistence diagram using Cubical Complex filtration can be achieved by $\Theta\left(3^{d}n+d2^{d}n\right)$, where $n$ is the number of vertices and $d$ is the dimension [11]. Then, the Wasserstein distance computation has a worst complexity of ${\mathcal{O}}(m^{2})$, where $m$ is the number of Birth-Death pairs in the persistence diagram. Both Topocode encoding and decoding include the computation of PDs and Wasserstein distance.
IV Simulation and Discussion